# Electrocardiogram–Language Model for Few-Shot Question Answering with Meta Learning

### v1

@@ Abstract
Electrocardiogram (ECG) interpretation requires specialized expertise, often involving synthesizing insights from ECG signals with complex clinical queries posed in natural language. The scarcity of labeled ECG data coupled with the diverse nature of clinical inquiries presents a significant challenge for developing robust and adaptable ECG diagnostic systems. This work introduces a novel multimodal meta-learning method for few-shot ECG question answering, addressing the challenge of limited labeled data while leveraging the rich knowledge encoded within large language models (LLMs). Our LLM-agnostic approach integrates a pre-trained ECG encoder with a frozen LLM (e.g., LLaMA and Gemma) via a trainable fusion module, enabling the language model to reason about ECG data and generate clinically meaningful answers. Extensive experiments demonstrate superior generalization to unseen diagnostic tasks compared to supervised baselines, achieving notable performance even with limited ECG leads. For instance, in a 5-way 5-shot setting, our method using LLaMA-3.1-8B achieves accuracy of 84.6%, 77.3%, and 69.6% on single verify, choose and query question types, respectively. These results highlight the potential of our method to enhance clinical ECG interpretation by combining signal processing with the nuanced language understanding capabilities of LLMs, particularly in data-constrained scenarios.
{IEEEkeywords}
electrocardiogram, few shot learning, large language models, meta learning, question answering, self-supervised learning
@@ 1 Introduction
[…]
Electrocardiograms (ECGs) provide a wealth of physiological information crucial for diagnosing a wide range of cardiac conditions. Interpreting these signals, however, requires specialized expertise, and even experienced clinicians can face challenges in discerning subtle patterns or adapting their knowledge to novel or complex cases. The advent of large language models (LLMs) coupled with advancements in multimodal learning presents a transformative opportunity to enhance ECG interpretation by integrating the rich contextual understanding of language with the detailed physiological insights encoded within ECG signals. This fusion of modalities allows for a more comprehensive and nuanced analysis, potentially leading to more accurate and timely diagnoses. Multimodal question answering (QA) systems, operating at this intersection of ECG data and natural language processing, are emerging as a powerful tool for automating and augmenting clinical workflows, offering the potential to improve diagnostic accuracy, efficiency, and accessibility. By enabling direct interaction with ECG data through natural language queries, these systems can streamline the diagnostic process and empower clinicians with more informed decision-making capabilities.
Developing robust and reliable multimodal QA systems for ECG interpretation relies on the availability of both high-quality and large quantities of labeled data. Yet, obtaining massive amounts of labeled ECGs from cardiologists is costly, which often results in limited datasets. Traditional supervised learning methods tend to perform well only on data with the same distribution as the training data. In real-world deployment, however, models frequently encounter new tasks and previously unseen populations outside the training distribution, where traditional methods may fail. Meta-learning [1, 2, 3], a paradigm focused on “learning to learn”, offers a compelling solution to this challenge. By training models on a diverse range of tasks, meta-learning enables them to acquire transferable knowledge and adapt rapidly to new, unseen tasks with minimal labeled data. This adaptive capacity is particularly valuable in the ECG-language QA domain, where new diagnostic questions and data distributions constantly emerge.
Figure 1: Illustration of class formation and attribute distribution for different question types.
Table 1: Overview of question types and data distribution within the restructured ECG-QA dataset used for few-shot learning.
Question Type Attributes Answers Classes (train:test) Example
Single-Verify 94 yes/no 156 (124:32)
@@
Single-Choose 165 both/none/attr_1/attr_2 250 (200:50)
@@
Single-Query 30 attr_1/attr_2/…/attr_n 260 (208:52)
@@
All 206 yes/no/both/none/…/attr_n 666 (532:134)
Few-shot learning (FSL), as a practical approach within meta-learning, shows significant promise in various medical imaging tasks [4]. The success of FSL underscores the potential of learning efficient representations that generalize effectively from limited examples [2]. While high-quality multimodal datasets, like those available for chest X-rays, have fueled progress in FSL for image-based diagnostics, the ECG domain lacks datasets specifically tailored for few-shot multimodal learning paradigms. The recent introduction of the ECG-QA dataset [5], built upon established ECG resources like PTB-XL [6] and MIMIC-IV-ECG [7], partially addresses this need with its diverse question types (single-verify, single-choose, single-query) and ECG attributes (e.g., SCP codes, noise types, heart axis deviations). However, this dataset lacks the structured task configurations necessary for developing and evaluating meta-learning models, leaving a significant gap in the advancement of ECG-language QA systems.
In response to these challenges, we propose a novel, LLM-agnostic, multimodal meta-learning framework specifically designed for few-shot ECG-language QA. Our architecture integrates a self-supervised pre-trained ECG encoder with a frozen LLM and a trainable multimodal fusion mapper bridging the ECG and language representations. This fusion mapper is crucial for acquiring transferable meta-knowledge, enabling rapid adaptation to new tasks. Furthermore, we curate a restructured variant of the ECG-QA dataset (see Figure 1 and Table 1), designed to facilitate meta-learning by creating diverse tasks with varying attribute-answer combinations. This restructuring allows us to rigorously evaluate a model’s ability to generalize to unseen diagnostic tasks in a few-shot setting. We demonstrate the effectiveness of our framework across a broad range of language models, showcasing superior generalization performance compared to fully supervised baselines in various few-shot settings and question types. Our findings highlight the potential of our approach to significantly impact clinical practice by enabling robust and adaptable ECG-language QA with limited labeled data.
@@ 2 Related Works
Deep learning has significantly advanced ECG interpretation, with models such as CNNs and Transformers demonstrating promising results in automated diagnosis [8, 9, 10]. However, these supervised approaches typically require large labeled datasets, hindering their generalizability to diverse patient populations and uncommon ECG presentations, a critical limitation in real-world clinical settings. While self-supervised learning methods offer a potential solution by learning from unlabeled ECG data [11, 12, 13, 14, 15], they have not yet been effectively leveraged for complex clinical question answering involving nuanced language understanding.
Multimodal learning has emerged as a powerful paradigm in healthcare, demonstrating success in integrating medical images with textual information [16, 17, 18, 19, 20, 21]. However, effectively fusing temporal physiological signals like ECG with the unstructured and often ambiguous nature of clinical language presents unique challenges, particularly in generative tasks like open-ended question answering. Our work directly addresses this gap by proposing a novel method for ECG-language fusion, enabling more comprehensive and nuanced diagnostic reasoning by leveraging the complementary information present in both modalities.
Furthermore, the inherent scarcity of labeled data for specific cardiac conditions necessitates efficient few-shot learning strategies. Meta-learning techniques, such as MAML [2], have shown promise in enabling rapid adaptation to new tasks with limited examples [22], offering a compelling approach for ECG interpretation. While recent studies have explored integrating LLMs with few-shot learning in medical domains [23, 24], the potential of combining meta-learning, LLMs, and multimodal fusion for ECG-language question answering remains largely unexplored. Our work contributes a method that integrates these key components, enabling adaptability to new tasks from limited labeled data while leveraging the powerful language understanding and generation capabilities of LLMs.
@@ 3 Methods
We present a method capable of rapidly adapting models to novel ECG question-answers (QAs) tasks with minimal labeled data. We frame this problem within the context of multimodal few-shot meta-learning. Here, we first define the meta-learning setting specific to ECG-language QAs, then detail the architecture of our proposed model, and finally describe the procedures for few-shot meta-training and inference.
@@ 3.1 Problem Formulation
We focus on the task of ECG-based question answering, where the goal is to predict an answer $a$ given an ECG signal $x$ and a natural language question $q$. Formally, we aim to learn a function $f_{\theta}$ parameterized by $\theta$: $a=f_{\theta}(x,q)$. Due to the scarcity of labeled data for certain ECG conditions and the need to generalize to emerging diseases, we adopt a few-shot learning approach. In this setting, we have access to a set of tasks, each consisting of a small support set and a query set. Specifically, let $\mathcal{D}_{\text{meta-train}}$ denote the meta-training dataset comprising $n$ tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{n}\}$, where each task $\mathcal{T}_{i}$ consists of a support set $D^{\text{s}}_{i}$ and a query set $D^{\text{q}}_{i}$: $\mathcal{T}_{i}=(D^{\text{s}}_{i},D^{\text{q}}_{i})$.
@@
Dataset Construction
We base our study on the Electrocardiogram Question Answering Dataset (ECG-QA) [5], which contains question-answer pairs annotated by expert clinicians and is built upon the PTB-XL [6] and MIMIC-IV-ECG [7] dataset. We focus on questions involving a single ECG and consider three types of questions:
@@
We prepare the dataset for few-shot meta learning by categorizing questions based on six types of attributes: SCP codes, noise types, stages of infarction, presence of ectopic beats, heart axis deviations, and numeric features. Each attribute encompasses multiple sub-attributes, leading to a diverse set of attribute-answer pairs. For instance, the SCP codes attribute includes specific diagnoses such as “non-diagnostic T-wave abnormalities“ and “conduction disturbances“.
Each class in our few-shot learning tasks corresponds to a unique attribute-answer pair. For the Single-Verify questions, classes are formed by pairs of attributes and binary answers (yes/no). Similarly, for Single-Choose questions, classes are based on attributes and possible options (both, none, specific sub-attributes), and for Single-Query questions, classes are defined by attributes and their specific values. Class formation and distribution in few-shot learning datasets for questions are shown in Figure 1.
We construct the meta-training dataset $\mathcal{D}_{\text{meta-train}}$ and the meta-testing dataset $\mathcal{D}_{\text{meta-test}}$ with mutually exclusive classes to evaluate the model’s ability to generalize to unseen attribute-answer pairs. Table 1 summarizes the number of attributes, possible answers, and classes for training and testing datasets in each question type. To ensure diversity and robustness, we include multiple question formulations with the same meaning but diverse expressions within each class. For example, the questions “Is non-diagnostic T-wave abnormality present in this ECG¿‘ and “Does this ECG reveal signs of non-diagnostic T-wave abnormalities¿‘ belong to the same class but provide variability in the language.
Figure 3: Illustration of the meta-training and meta-testing processes with Model-Agnostic Meta-Learning (MAML) [2].
Task Definition
@@ 3.2 Model Architecture
Our proposed model integrates ECG signal processing with natural language understanding to perform few-shot ECG-based question answering. The architecture consists of four main components: (1) a text tokenizer and embedder, (2) an ECG encoder, (3) a multimodal fusion module, and (4) a text decoder, as illustrated in Figure 2.
Text Encoder
@@
ECG Encoder
To extract meaningful representations from ECG signals, we pre-train an ECG encoder based on prior work [13]. Let $X=\{x_{1},x_{2},\ldots,x_{N}\}$ represent a set of ECG recordings, where each $x_{i}\in\mathbb{R}^{T_{s}\times C}$ corresponds to an ECG signal with $T_{s}$ time steps and $C$ leads. The ECG encoder processes each $x_{i}$ to produce a sequence of embeddings $E_{i}=\{e_{i,1},e_{i,2},\ldots,e_{i,K}\}$, capturing both local and global features of the ECG data.
The encoder incorporates techniques such as Wav2Vec (W2V), Contrastive Masked Segment Comparison (CMSC), and Random Lead Masking (RLM)[13], pre-trained on the PhysioNet 2021 dataset [25]. The W2V component uses convolutional and Transformer layers to derive contextualized representations from raw ECG signals. CMSC enhances temporal invariance by contrasting adjacent segments within ECG recordings. RLM improves generalization by masking random leads during training, enabling robustness across varying lead configurations.
Multimodal Fusion Mapper (Meta Mapper)
@@
Text Decoder (Language Model)
@@
3.3 Few-shot Meta Training and Inference
To enable rapid adaptation to new ECG question-answering tasks with minimal labeled data, we employ a few-shot meta-learning technique based on Model-Agnostic Meta-Learning (MAML)[2]. The meta-training process aims to find model parameters that are well-suited for quick fine-tuning on unseen tasks.
Meta-Training Phase
During meta-training, we sample a batch of tasks ${\mathcal{T}_{i}}$ from the task distribution $p(\mathcal{T})$. Each task ${\mathcal{T}_{i}}$ consists of a support set $D_{i}^{s}$ and a query set $D_{i}^{q}$. The support set contains $N$ classes with $K$ examples each ($N$-way $K$-shot learning), and the query set is used to evaluate adaptation performance.
@@ 3.3 Few-shot Meta Training and Inference
Outer Loop: Meta-Optimization
@@
Meta-Testing Phase

### v2

@@ Abstract
Electrocardiogram (ECG) interpretation requires specialized expertise, often involving synthesizing insights from ECG signals with complex clinical queries posed in natural language. The scarcity of labeled ECG data coupled with the diverse nature of clinical inquiries presents a significant challenge for developing robust and adaptable ECG diagnostic systems. This work introduces a novel multimodal meta-learning method for few-shot ECG question answering, addressing the challenge of limited labeled data while leveraging the rich knowledge encoded within large language models (LLMs). Our LLM-agnostic approach integrates a pre-trained ECG encoder with a frozen LLM (e.g., LLaMA and Gemma) via a trainable fusion module, enabling the language model to reason about ECG data and generate clinically meaningful answers. Extensive experiments demonstrate superior generalization to unseen diagnostic tasks compared to supervised baselines, achieving notable performance even with limited ECG leads. For instance, in a 5-way 5-shot setting, our method using LLaMA-3.1-8B achieves an accuracy of 84.6%, 77.3%, and 69.6% on single verify, choose and query question types, respectively. These results highlight the potential of our method to enhance clinical ECG interpretation by combining signal processing with the nuanced language understanding capabilities of LLMs, particularly in data-constrained scenarios.
Data and Code Availability This paper uses ECG-QA, the question-answering dataset for electrocardiogram (ECG) analysis (Oh et al., 2024) regarding PTB-XL (Wagner et al., 2020) and MIMIC-IV-ECG (Gow et al., 2023), separately available at https://github.com/Jwoo5/ecg-qa, https://physionet.org/content/ptb-xl/1.0.1/, https://physionet.org/content/mimic-iv-ecg/1.0/.
Institutional Review Board (IRB) Our research uses publicly available data, which does not require IRB approval.
@@ 1 Introduction
Electrocardiograms (ECGs) provide a wealth of physiological information crucial for diagnosing a wide range of cardiac conditions. Although doctors are professionally trained to diagnose (Garcia and Holtz, 2001; O’Keefe, 2008), and even AI systems have shown promise in not only enhancing diagnostic accuracy but also relieving the pressure on healthcare professionals (Jin et al., 2024; Ribeiro et al., 2020; Hannun et al., 2019). However, they are usually trained in limited and incomplete categories (Al-Alshaikh et al., 2024). The advent of large language models (LLMs) coupled with advancements in multimodal learning presents a transformative opportunity to enhance ECG interpretation by integrating the rich contextual understanding of language with the detailed physiological insights encoded within ECG signals. This fusion of modalities allows for a more comprehensive and nuanced analysis, potentially leading to more accurate and timely diagnoses. Multimodal question answering (QA) systems, operating at this intersection of ECG data and natural language processing, are emerging as a powerful tool for automating and augmenting clinical workflows, offering the potential to improve diagnostic accuracy, efficiency, and accessibility. By enabling direct interaction with ECG data through natural language queries, these systems can streamline the diagnostic process and empower clinicians with more informed decision-making capabilities.
Developing robust and reliable multimodal QA systems for ECG interpretation relies on the availability of both high-quality and large quantities of labeled data. Yet, obtaining massive amounts of labeled ECGs from cardiologists is costly, which often results in limited datasets. Traditional supervised learning methods tend to perform well only on data with the same distribution as the training data. In real-world deployment, however, models frequently encounter new tasks and previously unseen populations outside the training distribution, where traditional methods may fail. Meta-learning (Andrychowicz et al., 2016; Finn et al., 2017; Thrun and Pratt, 1998), a paradigm focused on “learning to learn”, offers a compelling solution to this challenge. By training models on a diverse range of tasks, meta-learning enables them to acquire transferable knowledge and adapt rapidly to new, unseen tasks with minimal labeled data. This adaptive capacity is particularly valuable in the ECG-language QA domain, where new diagnostic questions and data distributions constantly emerge.
Table 1: Overview of question types and data distribution within the meta learning benchmark dataset created for few-shot ECG question answering.
Question Type Attributes Answers Classes (train:test) Samples Example
Single-Verify 94 yes/no 156 (124:32) 34,105
@@
Single-Choose 165 both/none/attr_1/attr_2 262 (209:53) 47,655
@@
Single-Query 30 attr_1/attr_2/…/attr_n 260 (208:52) 63,125
@@
All 206 yes/no/both/none/…/attr_n 678 (541:137) 144,885
Few-shot learning (FSL), as a practical approach within meta-learning, shows significant promise in various medical imaging tasks (Pachetti and Colantonio, 2024). The success of FSL underscores the potential of learning efficient representations that generalize effectively from limited examples (Finn et al., 2017). While high-quality multimodal datasets, like those available for chest X-rays, have fueled progress in FSL for image-based diagnostics, the ECG domain lacks datasets specifically tailored for few-shot multimodal learning paradigms. The recent introduction of the ECG-QA dataset (Oh et al., 2024), built upon established ECG resources like PTB-XL (Wagner et al., 2020) and MIMIC-IV-ECG (Gow et al., 2023), partially addresses this need with its diverse question types (single-verify, single-choose, single-query) and ECG attributes (e.g., SCP codes, noise types, heart axis deviations). However, existing dataset lacks the structured task configurations necessary for developing and evaluating meta-learning models, leaving a significant gap in the advancement of ECG-language QA systems.
In response to these challenges, we propose a novel, LLM-agnostic, multimodal meta-learning framework specifically designed for few-shot ECG-language QA. Our architecture integrates a self-supervised pre-trained ECG encoder with a frozen LLM and a trainable multimodal fusion mapper bridging the ECG and language representations. This fusion mapper is crucial for acquiring transferable meta-knowledge, enabling rapid adaptation to new tasks. Furthermore, we create a benchmarking variant of the ECG-QA dataset (see Table 1), designed to facilitate meta-learning having diverse tasks with varying attribute-answer combinations. This benchmark dataset allows us to rigorously evaluate a model’s ability to generalize to unseen diagnostic tasks in a few-shot setting. We demonstrate the effectiveness of our framework across a broad range of language models, showcasing superior generalization performance compared to fully supervised baselines in various few-shot settings and question types. Our findings highlight the potential of our approach to significantly impact clinical practice by enabling robust and adaptable ECG-language QA with limited labeled data.
@@ 2 Related Works
Deep learning has significantly advanced ECG interpretation, with models such as CNNs and Transformers demonstrating promising results in automated diagnosis (Chugh and Jain, 2023; Woo et al., 2024; Sun et al., 2023). However, these supervised approaches typically require large labeled datasets, hindering their generalizability to diverse patient populations and uncommon ECG presentations, a critical limitation in real-world clinical settings. While self-supervised learning methods offer a potential solution by learning from unlabeled ECG data (Gopal et al., 2021; Tonekaboni et al., 2021; Oh et al., 2022; Saeed et al., 2019; Kiyasseh et al., 2021), they have not yet been effectively leveraged for complex clinical question answering involving nuanced language understanding.
Multimodal learning has emerged as a powerful paradigm in healthcare, demonstrating success in integrating medical images with textual information (Krones et al., 2025; Boecking et al., 2022; Zhang et al., 2020; Rasmy et al., 2021; Warner et al., 2024; Acosta et al., 2022). However, effectively fusing temporal physiological signals like ECG with the unstructured and often ambiguous nature of clinical language presents unique challenges, particularly in generative tasks like open-ended question answering. Our work directly addresses this gap by proposing a novel method for ECG-language fusion, enabling more comprehensive and nuanced diagnostic reasoning by leveraging the complementary information present in both modalities.
Furthermore, the inherent scarcity of labeled data for specific cardiac conditions necessitates efficient few-shot learning strategies. Meta-learning techniques, such as MAML (Finn et al., 2017), have shown promise in enabling rapid adaptation to new tasks with limited examples (Vettoruzzo et al., 2024), offering a compelling approach for ECG interpretation. While recent studies have explored integrating LLMs with few-shot learning in medical domains (Jin et al., 2023; Yu et al., 2023), the potential of combining meta-learning, LLMs, and multimodal fusion for ECG-language question answering remains largely unexplored. Our work contributes a method that integrates these key components, enabling adaptability to new tasks from limited labeled data while leveraging the powerful language understanding and generation capabilities of LLMs.
@@ 3 Methods
We present a method capable of rapidly adapting models to novel ECG Question-Answers (QAs) tasks with minimal labeled data. We frame this problem within the context of multimodal few-shot meta-learning consisting of three key phases. Here, we first define the meta-learning dataset specific to ECG-language QAs , where the objective is to classify unseen examples into one of N new ‘test’ classes, given only a few reference examples per class(Triantafillou et al., 2019). Then, we detail the architecture of our proposed model that integrates ECG analysis with question processing to generate the corresponding answer, and finally, we describe the procedures for few-shot meta-training and inference,in which a few gradient steps may provide strong results on a new task can be considered as constructing an internal representation that is generically applicable to numerous tasks.(Yuan and Nguyen, 2023)
@@ 3.1 Problem Formulation
We focus on the task of ECG-based question answering, where the goal is to predict an answer $a$ given an ECG signal $x$ and a natural language question $q$. Formally, we aim to learn a function $f_{\theta}$ parameterized by $\theta$: $a=f_{\theta}(x,q)$. Due to the scarcity of labeled data for certain ECG conditions and the need to generalize to emerging diseases, we adopt a few-shot learning approach. In this setting, we have access to a set of tasks, each consisting of a small support set and a query set. A single meta-learning step refers to an optimization after a support set (i.e., the few-shot samples) is used by the model to learn across different tasks and a query set adapts to a new task (Ravi and Larochelle, 2016). Specifically, let $\mathcal{D}_{\text{meta-train}}$ denote the meta-training dataset comprising $n$ tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{n}\}$, where each task $\mathcal{T}_{i}$ consists of a support set $D^{\text{s}}_{i}$ and a query set $D^{\text{q}}_{i}$: $\mathcal{T}_{i}=(D^{\text{s}}_{i},D^{\text{q}}_{i})$.
@@
Meta Learning Benchmark Dataset.
We create a benchmarking dataset for meta learning in our study using the ECG-QA dataset (Oh et al., 2024), which contains question-answer pairs annotated by expert clinicians and is built upon the PTB-XL (Wagner et al., 2020) and MIMIC-IV-ECG (Gow et al., 2023) datasets. We focus on questions involving a single ECG and consider three types of questions:
@@
We create our dataset for few-shot meta learning by categorizing questions based on six types of attributes: SCP codes, noise types, stages of infarction, presence of ectopic beats, heart axis deviations, and numeric features. Each attribute encompasses multiple sub-attributes, leading to a diverse set of attribute-answer pairs. For instance, the SCP codes attribute includes specific diagnoses such as “non-diagnostic T-wave abnormalities“ and “conduction disturbances“.
Each class in our few-shot learning tasks corresponds to a unique attribute-answer pair. For the Single-Verify questions, classes are formed by pairs of attributes and binary answers (yes/no). Similarly, for Single-Choose questions, classes are based on attributes and possible options (both, none, specific sub-attributes), and for Single-Query questions, classes are defined by attributes and their specific values.
We construct the meta-training dataset $\mathcal{D}_{\text{meta-train}}$ and the meta-testing dataset $\mathcal{D}_{\text{meta-test}}$ with mutually exclusive classes to evaluate the model’s ability to generalize to unseen attribute-answer pairs. Table 1 summarizes the number of attributes, possible answers, and classes for training and testing datasets in each question type. To ensure diversity and robustness, we include multiple question formulations with the same meaning but diverse expressions within each class. For example, the questions ”Is non-diagnostic T-wave abnormality present in this ECG?” and ”Does this ECG reveal signs of non-diagnostic T-wave abnormalities?” belong to the same class but provide variability in the language.
Task Definition.
@@ 3.2 Model Architecture
The architecture for ECG-based question answering consists of four main components: (1) a pretrained and frozen text tokenizer and embedder for semantic understanding of questions, (2) a pretrained and frozen ECG encoder for extracting meaningful representations from ECG signals, (3) a trainable multimodal fusion module to align ECG embeddings with the textual representation space, and (4) a text decoder to generate language-based answers, as illustrated in Figure 1.
Text Encoder.
@@
ECG Encoder.
To extract meaningful representations from ECG signals, we pre-train an ECG encoder based on prior work (Oh et al., 2022). Let $X=\{x_{1},x_{2},\ldots,x_{N}\}$ represent a set of ECG recordings, where each $x_{i}\in\mathbb{R}^{T_{s}\times C}$ corresponds to an ECG signal with $T_{s}$ time steps and $C$ leads. The ECG encoder processes each $x_{i}$ to produce a sequence of embeddings $E_{i}=\{e_{i,1},e_{i,2},\ldots,e_{i,K}\}$, capturing both local and global features of the ECG data.
The encoder incorporates techniques such as Wav2Vec (W2V), Contrastive Masked Segment Comparison (CMSC), and Random Lead Masking (RLM)(Oh et al., 2022), pre-trained on the PhysioNet 2021 dataset (Goldberger et al., 2000 (June 13). The W2V component uses convolutional and Transformer layers to derive contextualized representations from raw ECG signals. CMSC enhances temporal invariance by contrasting adjacent segments within ECG recordings. RLM improves generalization by masking random leads during training, enabling robustness across varying lead configurations.
Multimodal Fusion Mapper (Meta Mapper).
@@
Text Decoder (Language Model).
@@
3.3 Few-shot Meta Training and Inference.
To enable rapid adaptation to new ECG question-answering tasks with minimal labeled data, we employ a few-shot meta-learning technique based on Model-Agnostic Meta-Learning (MAML)(Finn et al., 2017). The meta-training process aims to find model parameters that are well-suited for quick fine-tuning on unseen tasks.
Meta-Training Phase.
During meta-training as shown in Appendix B.1 Figure 2, we sample a batch of tasks ${\mathcal{T}_{i}}$ from the task distribution $p(\mathcal{T})$. Each task ${\mathcal{T}_{i}}$ consists of a support set $D_{i}^{s}$ and a query set $D_{i}^{q}$. The support set contains $N$ classes with $K$ examples each ($N$-way $K$-shot learning), and the query set is used to evaluate adaptation performance.
@@
Outer Loop: Meta-Optimization.
@@
Meta-Testing Phase.